OpenStudy (anonymous):

Can someone please help me solve this one problem. I have a quiz tomorrow, and problems like these are the ones that are still confusing me: -Verify the following identities: cos(-theta)/ 1+sin(-theta) = sec(theta) + tan (theta)

OpenStudy (anonymous):

Also, if you know any pointers in tackling these problems, that will be greatly appreciated. :)

OpenStudy (amistre64):

it helps to define all the terms in sin cos functions

OpenStudy (amistre64):

cos/( 1+sin) = 1/cos + sin\cos cos/(1+sin) = (1+sin)/cos i dont think that is an identity. Identities are always equal.

OpenStudy (amistre64):

cos^2 = (1+sin)^2 cos^2 = 1+sin^2+2sin cos^2 - sin^2 = 1 + 2sin 1 - sin^2 - sin^2 = 1 + 2sin 1 -2sin^2 = 1 + 2sin -sin^2 = sin is not always true

OpenStudy (amistre64):

of course, i did not read the negative parts of the theta :/ oh well